2 L of an ideal gas at a pressure of 10 atm expands isothermally into a vacuum until its total volume is 10 L. How much work is done in the expansion?

Grade:12

1 Answers

Amit Sahoo
6 years ago
W=p(ext)×∆vW=1×(10-8). {P(ext)=p(atm)=1atm}W=8L-atm. {or 8×101.33=810.64 Joule}Here pressure given is not the external pressure and hence pressure acting is only the atmospheric pressure and I.e. 1atm....